# Inflated hot Jupiters: inferring average atmospheric velocity via Ohmic models coupled with internal dynamo evolution

### v1

@@ 3.1 Structure equations
We model the evolution of HJs using the publicly available code MESA (Paxton et al., 2011, 2013, 2019), and solving the one dimensional time-dependent equations of internal structure applied to gas giants. The first is the conservation of mass,
@@
where $m$ is the mass enclosed within the radius $r$, and $\rho$ is the density. The second is the hydrostatic equilibrium,
@@
where $p$ is the pressure and $G$ the gravitational constant. The energy transport equation is given by
@@
where $T$ is the temperature and $\nabla=d\ln T/d\ln p$ is the logarithmic temperature gradient, set to the smallest between the adiabatic gradient and the radiative gradient.
The equations above define the internal structure at a given time; its evolution is driven by energy conservation, which in terms of the internal luminosity $L$ reads
$\frac{dL}{dm}=\epsilon_{\rm grav}+\epsilon_{\rm irr}+\epsilon_{j}\,,$ (5)
where $\epsilon_{\rm grav}=-TdS/dt$ is the heating term due to gravitational contraction or inflation, with $S$ being the specific entropy; $\epsilon_{\rm irr}$ is heating of the outermost layers due to irradiation (see Sect. 3.3), while $\epsilon_{j}$ represents any additional heat deposition rate, which in our case comes from Ohmic dissipation of currents and varies in depth and time, as specified in Sect. 4. The planetary radius $R$ varies in time, according to these structure equations. MESA typically solves the equations from the centre to the outermost layers (located at $p\sim 0.1$ bar).
@@
An important free parameter in solving the structure is the total planetary mass $M=m(R)$, which is assumed to be constant throughout the evolution. We do not consider any mass loss due to, for example, wind stripping of the atmosphere (see e.g. Lazovik 2023). Such processes could explain at least part of the HJs with $R\lesssim 1\leavevmode\nobreak\ R_{j}$ and have been inferred in some short-orbit exoplanets, via Lyman-$\alpha$ and helium triplet observations (e.g., Salz et al. 2015; Lavie et al. 2017; Tyler et al. 2024). For HJs, observations indicate a relatively small mass loss ($\lesssim 1\%$) accumulated over their lifetime. However, this might be an observational bias, and HJ evaporation could affect the observed mass distribution (Kurokawa & Nakamoto, 2014), similar to the atmospheric loss observed in hot super-Earths, which can leave behind smaller, gas-depleted rocky remnants (e.g., Owen & Wu 2016). These effects should be more important for relatively light HJs and hot Saturns (Thorngren et al., 2023). Therefore, this study mostly focuses on the upper range of the radius values.
@@ 3.2 Core mass
In this paper, we fix the core mass to $M_{c}=10\sqrt{M/M_{j}}\leavevmode\nobreak\ M_{\oplus}$, compatible with the trend inferred from the sample of cold Jupiters by Thorngren et al. (2016). In order to fix the core density $\rho_{c}$, we consider what is the central density obtained at 10 Gyr with a model having the same total mass $M$, but in absence of both core and irradiation (i.e., the largest value for central density, since it increases with age and decreases with irradiation and age). In our models, such a central density goes approximately as $\sim 2.7(M/M_{j})$  g cm-3  between 0.3 and 5 $M_{j}$ (becomes slightly steeper for higher masses). Therefore, we conservatively fix a core density $\rho_{c}=10(M/M_{j})$  g cm-3, which recovers values often fixed in previous works using MESA for a $1\leavevmode\nobreak\ M_{j}$ planet, e.g. Komacek & Youdin (2017). This practical prescription ensures that the central density, $\rho_{c}$, never falls below the density of the deepest envelope layers at any time ($t\leq 10$ Gyr) or for any model parameters ($M$ and $T_{\rm eq}$), thereby avoiding unphysical values. Moreover, these values are well below what can be extrapolated by the iron equation of state (the most dense case, e.g. Smith et al. 2018), meaning that the core here considered is compatible with representing material less dense than iron. In any case, for the core masses here considered, the role of $\rho_{c}$ is minor, since it is only used to fix the innermost values of enclosed mass, $m_{i}=M_{c}$, and radius, $r_{i}=[3M_{c}/(4\pi\rho_{c})]^{1/3}$.
In App. C we briefly validate standard cooling results and show the sensitivity of results (without internal heating) for planets of different core mass, core density, and composition. Since these are well-known results and are not the objective of the paper, hereafter we will vary only the observables, $M$ and $T_{\rm eq}$, fixing the composition to Solar values, and the values of core mass and density as described above.
@@ 3.3 Irradiation
MESA provides a few different ways to implement an irradiated atmosphere. As in Komacek & Youdin (2017), we use a specific energy absorption rate, $\epsilon_{\rm irr}=F_{\rm irr}/(4\Sigma_{\star})$, applied uniformly through the outermost mass column $m(r)\leqslant\Sigma_{\star}$, where $F_{\rm irr}=4\sigma_{\rm sb}T_{\rm eq}^{4}$ is the irradiation flux, $\sigma_{\rm sb}$ is the Stefan-Boltzmann constant, and $\Sigma_{\star}$ is the heated column mass, a value which parametrizes the atmospheric absorption. We consider a constant value of stellar flux $F_{\rm irr}$, thus neglecting its increase during the stellar evolution. This effect has been considered instead in e.g. Komacek et al. (2020) and Lopez & Fortney (2016), who find that planetary reinflation is possible (assuming a constant Ohmic efficiency), especially for $\lesssim 1\leavevmode\nobreak\ M_{j}$ planets around stars at the end of the main sequence phase. Observational evidence in this sense might have been found (Hartman et al., 2016).
@@
A more detailed atmospheric modelling, accounting quantitatively for chemical composition and opacity, would give a different, non-uniform and non-grey heat deposition profile $\epsilon_{\rm irr}$. Note also that, since the model is 1D, we cannot model the differences of irradiation across the surface. Hence, the local values of the temperature in the outer layer must be interpreted as averages. While these caveats would be important if one wants to fit, e.g., the atmospheric properties of a given object, this paper is focused on the deep deposition of heat and aims at an exploration of the main parameters to explain the trends in the entire sample. Therefore, for our purposes, this simple implementation is sufficient.
@@
We fix the absorbing mass column to $\Sigma_{\star}=250$  g cm-2, which approximately corresponds to a visible, grey opacity $\kappa_{\rm vis}=4\times 10^{-3}$ cm2 g-1 (as in previous similar works, e.g. Fortney et al. 2008; Guillot 2010; Owen & Wu 2016; Komacek & Youdin 2017). This value, for an $M=1M_{j}$ planet at Gyr ages, corresponds to absorbing stellar energy homogeneously (per unit mass) down to a depth of $p\sim 1$ bar. We show in App. C that the results depend only slightly on the values of $\Sigma_{\star}$ within a reasonable range, with planetary radius variations of a few percent at most, much less than the main parameters of the model ($M$, $T_{\rm eq}$, and the Ohmic parameters introduced below).
@@
Practically, we consider the previously built planetary structure (fixing the mass, core mass and density, composition) and find a relaxed solution considering irradiation. Such a profile is the starting point for the evolution. This extra step helps the convergence at early times, especially for high values of irradiation and low masses. We start the simulations with an initial radius of $\sim$4 $R_{j}$, a value close to what chosen in other works (Paxton et al., 2013), and much larger than the evolved values at the relevant ages, so that results at Gyr are insensitive to the specific choice.
@@ 4 Ohmic heating model
$\epsilon_{j}=\frac{Q_{j}}{\rho}=\frac{J^{2}}{\sigma\leavevmode\nobreak\ \rho}\,,$ (6)
where $Q_{j}=J^{2}/\sigma$ is the volumetric Ohmic heating rate, $J$ is the electrical current density and $\sigma$ the electrical conductivity. Obviously, the radial profiles of these quantities represent the averages over each sphere $r$. The Ohmic model we describe hereafter is implemented as a modification of the MESA make_planet suite (Paxton et al., 2013), in which we define the extra heat source $\epsilon_{j}$, with all the ingredients and calculations explained in the following subsections. The module is publicly available [^3] https://github.com/danielevigano/mesa_ohmic_hj and has been developed for MESA v24.08.1, the version we use in this work.
@@ 4.1 Electrical conductivity
At shallow layers of HJs ($\rho\lesssim 10^{-2}$ g cm-3, i.e. $p\lesssim$ kbar for Gyr ages), the electrical conductivity is dominated by the thermal ionization of the alkali metals (e.g., Kumar et al. 2021). As the density increases ($\rho\gtrsim 10^{-2}$  g cm-3), the electrons start becoming degenerate and ions becoming strongly coupled (e.g., Redmer 1997; Ramakrishna et al. 2024), so that the pressure-ionization from hydrogen and helium (the most abundant species), negligible at low densities, produces a sudden increase of conductivity with density, which is more drastic at lower temperatures. Pressure ionization is a many-particle effect that arises due to the lowering of ionization energy in a high-density plasma. When degeneracy and strong correlations plays a role, more detailed calculations, usually based on density functional theory and molecular dynamics (DFT-MD), are needed. The recent review by Bonitz et al. (2024) (see also the extensive list of references within) gives an overview on the transport properties calculations for such warm dense matter at high densities, $\rho\gtrsim 10^{-2}$  g cm-3.
@@
Above $\rho\gtrsim 1$  g cm-3, the conductivity has a reduced dependence on the temperature and follows the trend shown for the Jovian adiabat by French et al. (2012), with precise theoretical values typically differing by a factor of a few from one model to another. On the other hand, at intermediate depths $\rho\sim 10^{-2}-1$  g cm-3, the pressure ionization contribution to the conductivity is highly sensitive on temperature, and the available points in the $\rho$-$T$ plane are few and especially sparse at $T\lesssim 10^{4}$ K, which is the regime of our interest. In this range, the conductivity changes abruptly with temperature (at a given $\rho$) and becomes very steep, see e.g. Holst et al. 2011.
@@
To incorporate all these results and ensure a correct transition from thermally-dominated to pressure-dominated ionization, we linearly interpolate444 We interpolate in the logarithm of the variables, as they span several orders of magnitude and are strictly positive. Due to the scarcity and uncertainty of available data across broad regions of the $\rho$$T$ parameter space of interest higher-order interpolations are not warranted. the corresponding tables for $\sigma$ in the $\log(\rho)$-$\log(T)$ plane: the ones from Kumar et al. (2021) (for $\rho\leq 10^{-2}$  g cm-3) and Fig. 42 of Bonitz et al. (2024) (for $\rho\geq 10^{-1}$  g cm-3). In the intermediate region ($\rho\sim 10^{-2}$-$10^{-1}$  g cm-3), we interpolate between the closest extremes of the two tables, again linearly with logarithmic quantities. Note that the latest DFT-MD results differs at deep layers from the values of $\sigma$ used in the HJ inflation study by e.g. Huang & Cumming (2012) and Wu & Lithwick (2013), who considered the thermal and pressure ionization via the electron fraction from the Saumon et al. (1995) equation of state, but employed the electron–neutral collision cross-section from Draine et al. (1983) and an analytical expression for electron-proton scattering. The expression used for electron-proton scattering in their conductivity calculations may lack accuracy over the temperature and density ranges relevant to Ohmic dissipation studies. Furthermore, their approach neglects the influence of strong coupling effects on electrical conductivity.
@@
Fig. 2 shows the conductivity profiles obtained at three ages ($t\sim 0.4,1.1,5$ Gyr) for four different models: $T_{\rm eq}=1500,2000$ K, and with (blue and red lines) or without (black and grey lines) the inclusion of Ohmic heating (see below). While in the external part the values are highly sensitive to the local temperature, which is higher for highly irradiated and/or internally heated cases, at $p\gtrsim 10^{5}$ bar the curves tend to similar, very high values of $\sigma$, as in French et al. (2012). In the same plot, we also show for comparison (brown, gold lines), the widely used (Perna et al., 2010a, b; Soriano-Guerrero et al., 2023) analytical approximation for the thermal ionization of potassium (the dominant one in the outermost regions of the envelope) neglecting degeneracy and strong coupling effects, and valid for not too high temperatures, $T\lesssim 2000$ K (see discussion in Soriano-Guerrero et al. 2023). It can be seen how substantial differences are encountered at $p\gtrsim 1000$ bar.
@@
Finally, note that the dissociation of molecular into atomic hydrogen is taken into account in the conductivity calculations (Kumar et al., 2021) here used, but not as a source of heat, which might become relevant in the uppermost layers for $T_{\rm eq}\gtrsim 2000$ K (Bell & Cowan, 2018; Gandhi et al., 2020). At those shallow layers, in any case, there a number of non-ideal MHD effects can take place (Koskinen et al., 2014), but the heat dissipation doesn’t affect the radius inflation, which is much more sensitive on deep deposition.

### v2

@@ 3.1 Structure equations
We modelled the evolution of HJs using the publicly available code MESA (Paxton et al., 2011, 2013, 2019) and solving the one-dimensional time-dependent equations of internal structure applied to gas giants. The first equation is the conservation of mass,
@@
where $m$ is the mass enclosed within the radius, $r$, and $\rho$ is the density. The second is the hydrostatic equilibrium,
@@
where $p$ is the pressure and $G$ is the gravitational constant. The energy transport equation is given by
@@
where $T$ is the temperature and $\nabla=d\ln T/d\ln p$ is the logarithmic temperature gradient, set to the smallest value between the adiabatic and the radiative gradients.
The equations above define the internal structure at a given time; its evolution is driven by energy conservation, which in terms of the internal luminosity, $L$, reads as
$\frac{dL}{dm}=\epsilon_{\rm grav}+\epsilon_{\rm irr}+\epsilon_{\rm j}\,,$ (5)
where $\epsilon_{\rm grav}=-TdS/dt$ is the heating term due to gravitational contraction or inflation, with $S$ as the specific entropy; $\epsilon_{\rm irr}$ is the heating of the outermost layers due to irradiation (see Sect. 3.3); and $\epsilon_{\rm j}$ represents any additional heat deposition rate, which in our case comes from Ohmic dissipation of currents and varies in depth and time, as specified in Sect. 4. The planetary radius $R$ varies in time, according to these structure equations. MESA typically solves the equations from the centre to the outermost layers (located at $p\sim 0.1$ bar).
@@
An important free parameter in solving the structure is the total planetary mass $M=m(R)$, which is assumed to be constant throughout the evolution. We do not consider any mass loss due to, for example, wind stripping of the atmosphere (see e.g. Lazovik 2023). Such processes could explain at least part of the HJs with $R\lesssim 1\leavevmode\nobreak\ R_{\rm j}$ and have been inferred in some short-orbit exoplanets, via Lyman-$\alpha$ and helium triplet observations (e.g. Salz et al. 2015; Lavie et al. 2017; Tyler et al. 2024). For HJs, observations indicate a relatively small mass loss ($\lesssim 1\%$) accumulated over their lifetime. However, this might be an observational bias, and HJ evaporation could affect the observed mass distribution (Kurokawa & Nakamoto, 2014), similar to the atmospheric loss observed in hot super-Earths, which can leave behind smaller, gas-depleted rocky remnants (e.g. Owen & Wu 2016). These effects should be more important for relatively light HJs and hot Saturns (Thorngren et al., 2023). Therefore, this study mostly focuses on the upper range of the radius values.
@@ 3.2 Core mass
In this paper, we fix the core mass to $M_{c}=10\sqrt{M/M_{\rm j}}\leavevmode\nobreak\ M_{\oplus}$, compatible with the trend inferred from the sample of cold Jupiters by Thorngren et al. (2016). In order to fix the core density $\rho_{c}$, we consider what is the central density obtained at 10 Gyr with a model having the same total mass $M$, but in absence of both core and irradiation (i.e. the largest value for central density, since it increases with age and decreases with irradiation and age). In our models, such a central density goes approximately as $\sim 2.7(M/M_{\rm j})$  g cm-3  between 0.3 and 5 $M_{\rm j}$ (becomes slightly steeper for higher masses). Therefore, we conservatively fix a core density $\rho_{c}=10(M/M_{\rm j})$  g cm-3, which recovers values often fixed in previous works using MESA for a $1\leavevmode\nobreak\ M_{\rm j}$ planet, e.g. Komacek & Youdin (2017). This practical prescription ensures that the central density, $\rho_{c}$, never falls below the density of the deepest envelope layers at any time ($t\leq 10$ Gyr) or for any model parameters ($M$ and $T_{\rm eq}$), thereby avoiding unphysical values. Moreover, these values are well below what can be extrapolated by the iron equation of state (the most dense case, e.g. Smith et al. 2018), meaning that the core here considered is compatible with representing material less dense than iron. In any case, for the core masses here considered, the role of $\rho_{c}$ is minor, since it is only used to fix the innermost values of enclosed mass, $m_{i}=M_{c}$, and radius, $r_{i}=[3M_{c}/(4\pi\rho_{c})]^{1/3}$.
In App. C we briefly validate standard cooling results and show the sensitivity of results (without internal heating) for planets of different core mass, core density, and composition. Since these are well-known results and are not the objective of the paper, hereafter we vary only the observables, $M$ and $T_{\rm eq}$, fixing the composition to Solar values, and the values of core mass and density as described above.
@@ 3.3 Irradiation
MESA provides a few different ways to implement an irradiated atmosphere. As in Komacek & Youdin (2017), we use a specific energy absorption rate, $\epsilon_{\rm irr}=F_{\rm irr}/(4\Sigma_{\star})$, applied uniformly through the outermost mass column $m(r)\leqslant\Sigma_{\star}$, where $F_{\rm irr}=4\sigma_{\rm sb}T_{\rm eq}^{4}$ is the irradiation flux, $\sigma_{\rm sb}$ is the Stefan-Boltzmann constant, and $\Sigma_{\star}$ is the heated column mass, a value which parametrises the atmospheric absorption. We consider a constant value of stellar flux $F_{\rm irr}$, thus neglecting its increase during the stellar evolution. This effect has been considered instead in e.g. Komacek et al. (2020) and Lopez & Fortney (2016), who find that planetary reinflation is possible (assuming a constant Ohmic efficiency), especially for $\lesssim 1\leavevmode\nobreak\ M_{\rm j}$ planets around stars at the end of the main sequence phase. Observational evidence in this sense might have been found (Hartman et al., 2016).
@@
A more detailed atmospheric modelling, accounting quantitatively for chemical composition and opacity, would give a different, non-uniform and non-grey heat deposition profile $\epsilon_{\rm irr}$. We also note that since the model is 1D, we cannot model the differences of irradiation across the surface. Hence, the local values of the temperature in the outer layer must be interpreted as averages. While these caveats would be important if one wants to fit, for example, the atmospheric properties of a given object, this paper is focused on the deep deposition of heat and aims at an exploration of the main parameters to explain the trends in the entire sample. Therefore, for our purposes, this simple implementation is sufficient.
@@
We fixed the absorbing mass column to $\Sigma_{\star}=250$  g cm-2, which approximately corresponds to a visible grey opacity $\kappa_{\rm vis}=4\times 10^{-3}$ cm2 g-1 (as in previous similar works, e.g. Fortney et al. 2008; Guillot 2010; Owen & Wu 2016; Komacek & Youdin 2017). This value, for an $M=1M_{\rm j}$ planet at gigayear ages, corresponds to absorbing stellar energy homogeneously (per unit mass) down to a depth of $p\sim 1$ bar. We show in App. C that the results depend only slightly on the values of $\Sigma_{\star}$ within a reasonable range, with planetary radius variations of a few percent at most, much less than the main parameters of the model ($M$, $T_{\rm eq}$, and the Ohmic parameters introduced below).
@@
Practically, we considered the previously built planetary structure (fixing the mass, core mass and density, composition) and found a relaxed solution considering irradiation. Such a profile is the starting point for the evolution. This extra step helps the convergence at early times, especially for high values of irradiation and low masses. We started the simulations with an initial radius of $\sim$4 $R_{\rm j}$, a value that is close to what was chosen in other works (Paxton et al., 2013) and much larger than the evolved values at the relevant ages so that the results at gigayears are insensitive to the specific choice.
@@ 4 Ohmic heating model
$\epsilon_{\rm j}=\frac{Q_{\rm j}}{\rho}=\frac{J^{2}}{\sigma\leavevmode\nobreak\ \rho}\,,$ (6)
where $Q_{\rm j}=J^{2}/\sigma$ is the volumetric Ohmic heating rate, $J$ is the electrical current density, and $\sigma$ is the electrical conductivity. Obviously, the radial profiles of these quantities represent the averages over each sphere, $r$. The Ohmic model we describe hereafter is implemented as a modification of the MESA make_planet suite (Paxton et al., 2013), in which we define the extra heat source, $\epsilon_{\rm j}$, with all the ingredients and calculations explained in the following subsections. The module is available [^3] https://github.com/danielevigano/mesa_ohmic_hj and has been developed for MESA v24.08.1, the version we use in this work.
@@ 4.1 Electrical conductivity
Within shallow layers of HJs ($\rho\lesssim 10^{-2}$ g cm-3, i.e. $p\lesssim$ kbar for Gyr ages), the electrical conductivity is dominated by the thermal ionisation of the alkali metals (e.g. Kumar et al. 2021). As the density increases ($\rho\gtrsim 10^{-2}$  g cm-3), the electrons start becoming degenerate and ions becoming strongly coupled (e.g. Redmer 1997; Ramakrishna et al. 2024), so that the pressure-ionisation from hydrogen and helium (the most abundant species), negligible at low densities, produces a sudden increase of conductivity with density, which is more drastic at lower temperatures. Pressure ionisation is a many-particle effect that arises due to the lowering of ionisation energy in a high-density plasma. When degeneracy and strong correlations plays a role, more detailed calculations, usually based on density functional theory and molecular dynamics (DFT-MD), are needed. The recent review by Bonitz et al. (2024) (see also the extensive list of references within) gives an overview on the transport properties calculations for such warm dense matter at high densities, $\rho\gtrsim 10^{-2}$  g cm-3.
@@
Above $\rho\gtrsim 1$  g cm-3, the conductivity has a reduced dependence on the temperature and follows the trend shown for the Jovian adiabat by French et al. (2012), with precise theoretical values typically differing by a factor of a few from one model to another. On the other hand, at intermediate depths $\rho\sim 10^{-2}-1$  g cm-3, the pressure ionisation contribution to the conductivity is highly sensitive on temperature, and the available points in the $\rho$-$T$ plane are few and especially sparse at $T\lesssim 10^{4}$ K, which is the regime of our interest. In this range, the conductivity changes abruptly with temperature (at a given $\rho$) and becomes very steep, see e.g. Holst et al. 2011.
@@
To incorporate all these results and ensure a correct transition from thermally dominated to pressure-dominated ionisation, we linearly interpolated444 We interpolate in the logarithm of the variables, as they span several orders of magnitude and are strictly positive. Due to the scarcity and uncertainty of available data across broad regions of the $\rho$$T$ parameter space of interest higher-order interpolations are not warranted. the corresponding tables for $\sigma$ in the $\log(\rho)$-$\log(T)$ plane: the conductivity values from Kumar et al. (2021) (for $\rho\leq 10^{-2}$  g cm-3) and Fig. 42 of Bonitz et al. (2024) (for $\rho\geq 10^{-1}$  g cm-3). In the intermediate region ($\rho\sim 10^{-2}$-$10^{-1}$  g cm-3), we interpolated between the closest extremes of the two tables, again linearly with logarithmic quantities. We note that the latest DFT-MD results differs at deep layers from the values of $\sigma$ used in the HJ inflation study by e.g. Huang & Cumming (2012) and Wu & Lithwick (2013), who considered the thermal and pressure ionisation via the electron fraction from the Saumon et al. (1995) equation of state but employed the electron–neutral collision cross-section from Draine et al. (1983) and an analytical expression for electron-proton scattering. The expression used for electron-proton scattering in their conductivity calculations may lack accuracy over the temperature and density ranges relevant to Ohmic dissipation studies. Furthermore, their approach neglects the influence of strong coupling effects on electrical conductivity.
@@
Figure 2 shows the conductivity profiles obtained at three ages ($t\sim 0.4,1.1,5$ Gyr) for four different models: $T_{\rm eq}=1500,2000$ K, and with (blue and red lines) or without (black and grey lines) the inclusion of Ohmic heating (see below). While in the external part the values are highly sensitive to the local temperature, which is higher for highly irradiated and/or internally heated cases, at $p\gtrsim 10^{5}$ bar the curves tend to similar, very high values of $\sigma$, as in French et al. (2012). In the same plot, we also show for comparison (brown, gold lines), the widely used (Perna et al., 2010a, b; Soriano-Guerrero et al., 2023) analytical approximation for the thermal ionisation of potassium (the dominant one in the outermost regions of the envelope) neglecting degeneracy and strong coupling effects, and valid for not too high temperatures, $T\lesssim 2000$ K (see discussion in Soriano-Guerrero et al. 2023). It can be seen how substantial differences are encountered at $p\gtrsim 1000$ bar.
@@
Finally, we note that the dissociation of molecular into atomic hydrogen is taken into account in the conductivity calculations (Kumar et al., 2021) here used, but not as a source of heat, which might become relevant in the uppermost layers for $T_{\rm eq}\gtrsim 2000$ K (Bell & Cowan, 2018; Gandhi et al., 2020). At those shallow layers, in any case, there a number of non-ideal MHD effects can take place (Koskinen et al., 2014), but the heat dissipation does not affect the radius inflation, which is much more sensitive on deep deposition.